Help me Please!!
Solve the following system of equations:
5X + 8Y = 1 and 3X – Y = -8

Answer by checkley75(3666)

5X+8Y=1
3X-Y=-8 MULTIPLY THIS EQUATION BY 8 & ADD.
24X-8Y=-64
5X+8Y=1
---------------------------------------------
29X=-64+1
29X=-63
X=-63/29 ANSWER.
5*-63/29+8Y=1
-315/29+8Y=1
8Y=1+315/29
8Y=(29+315)/29
8Y=344/29
Y=344/29*1/8
Y=43/29 ANSWER.
PROOF
3*-63/29-43/29=-8
-189/29-43/29=-8
(-189-43)/29=-8
-232/29=-8
-8=-8